If the product water is liquid, what is the open circuit voltage of a H2 fuel cell if it operates at 70oC with pH2=2atm, pO2=2atm. Do not consider the overvoltages.
Answer #1

Hi
The reaction for the standard Hydrogen reaction is -
H2+0.5O2=H20
The value of the standard potential 1.23 V.

From Nerst equation
E=Eo+(RT/nF)*lnK

We know that K is the value of Kp ,since water is in aqueous state. We can take its activity to be 1.

E=1.23+(8.314*343/2*96500)*ln(xH2*xO2^0.5)

Total pressure can be taken as the sum of the pressure of H2 and O2 because H20 is in the form of water.
xH2 = 0.5
xO2 = 0.5


E=1.23+(8.314*343/2*96500)*ln(0.5*0.5^0.5)

E=1.21 V

Hope this helps.

Feel free to comment in the case of any doubt.
Thank you